# Incremental Generation is Necessity and Sufficient for Universality in Flow-Based Modelling

### v1

@@ 5 Conclusion
In this article, we have shown that incremental generation is both necessary and sufficient for universal flow-based approximation, while non-incremental flows form a meagre, non-universal subset. By coupling dynamical, algebraic, and approximation-theoretic tools, we established quantitative rates and demonstrated that lifted incremental flows yield structured universality for both functions and probability measures.
@@
The authors would like to thank Paulo Varandas for his encouraging suggestion for proving the meagerness of the set of flows and to Giulia Livieri for her incredibly helpful feedback during the final stages of the manuscript. A. Kratsios and H. Rouhvarzi would like to acknowledge financial support from an NSERC Discovery Grant No. RGPIN-2023-04482 and No. DGECR-2023-00230. They also acknowledge that resources used in preparing this research were provided, in part, by the Province of Ontario, the Government of Canada through CIFAR, and companies sponsoring the Vector Institute [^6] https://vectorinstitute.ai/partnerships/current-partners/.
@@ Appendix A Proof of the Negative Result
Then, given any neighborhood $U$ of the image of $\gamma$, there exists a flow $\psi$ on $M$ satisfying:
@@
Also, Lemma 16 property (a) makes the perturbed function in Lemma 17 to have support on a finite number of neighborhoods of the curves joining $p_{i}$ to $q_{i}$. As $h$ is the result of Lemma 17 take $\eta$ and those neighborhoods small enough such that these neighborhoods stay inside $P$. This makes the function $h$ to be compactly supported on $P$. As a result, $g$ must be compactly supported on $P$.
@@
As $D\text{exp}_{x}(0)=Id$ there exists $L>1$ tending to 1 as $G\rightarrow 0$ s.t. :
$d(f_{1}(z),f(z))\leq L\|(F-R)\circ\mathrm{exp}_{x}^{-1}(z)\|\leq L\|F-R\|G\leq L\|F-R\|\varepsilon$ (16)
@@
For every $C^{0}$-Perturbation $g$ of $f_{2}$ in $\operatorname{\mathcal{H}}_{d}(P)$ one has: $0\neq\text{deg}(f_{2}^{k}(x)-x,0)=\text{deg}(g^{k}(x)-x,0)$. There exists a point $z$ s.t. $g^{k}(z)-z=0$ so $g^{k}$ has a fixed point. Thus, $g$ has a periodic point of period $k>1$ that is not contained in a continuum subset of the set of $k$-periodic points. It means $f_{2}$ has an open neighborhood. This proves the openness.
Proof of Theorem 1: Since every neural ODE is, by definition, a flowable homeomorphisms supported on any such $\delta,1-\delta]^{d}$ then our next results necessitates that the set of neural ODEs is itself meager in $\operatorname{\mathcal{H}}_{d}([0,1]^{d})$ as well.
@@ B.1.1 General Case
Proof of Proposition 6: Consider vector fields $V^{\prime}(x,y)=(-\pi(y-\frac{1}{2}),\pi(x-\frac{1}{2}))$ and $W^{\prime}(x,y)=(-x+\frac{1}{2},0)$ which are $\pi$ rotation around the point $p=(\frac{1}{2},\frac{1}{2})$ and squeeze to the line $x=\frac{1}{2}$ (call this line $l$). Consider the bump function $\rho$ which is $1$ on the circle with radius $r=\frac{1}{8}$ around $p$ and zero out of the circle with radius $R=\frac{1}{4}$. $V=V^{\prime}\circ\rho,W=W^{\prime}\circ\rho$ are compactly supported on $0,1]^{d}$. Denote time one map of $V,W$ as $\varphi_{V},\varphi_{W}$. Then $\varphi_{V}\circ\varphi_{W}\notin\mathrm{Flow}$ . suppose not, i.e. there exists a vector field $F$ which its time one flow $\varphi_{F}$ is equal to $\varphi_{V}\circ\varphi_{W}$.
Note that for the composition $\varphi_{V}\circ\varphi_{W}$ any point $a\in B_{r}(p)\setminus l$ converges to zero on iteration but any $b\in l$ has periodic point two. Consider the trajectory $\gamma$ of the flow of $\varphi_{V}\circ\varphi_{W}$. For any point $c=r_{c}e^{i\theta}$ (considering polar coordinate) on a neighborhood $S$ of the point $\varphi_{V}\circ\varphi_{W}(b)$ on $\gamma$, radius of $\varphi_{V}\circ\varphi_{W}(c)$ will be smaller than $r_{c}$ as a result radius of any point other than $b$ on $(\varphi_{V}\circ\varphi_{W})^{2}(S)$ will be smaller but $(\varphi_{V}\circ\varphi_{W})^{2}(b)=b$ so there will be a discontinuity at the point $b$ but all periodic orbits of a continuous flow form a continuum (i.e. a compact and connected metric space that contains at least two points) which is a contradiction.
@@
Proof of  Theorem 7: As $\varphi\in\operatorname{\mathcal{H}}_{d}^{G,s}([0,1]^{d})$ in (7) define $\varphi_{g}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\operatorname{Flow}{}(V_{g})$ for every $g\in[G]$. So one can write:
$\varphi=\bigcirc_{g=1}^{G}\,\varphi^{(g)}(x)$ (20)
We use Lemma 19 for each $\varphi^{(g)}$ to find Neural ODEs $\{\Psi^{(g)}\}_{g=1}^{G}$. One can write $\Psi^{(g)}(x)=z_{1}^{x}$ as in (17), then as $\Phi^{(g)}$ is Lipschitz we can write:
$\big\|z_{t}^{x}-z_{t}^{y}\big\|_{l^{\infty}}\leq\big\|x-y\big\|_{l^{\infty}}+\int_{0}^{t}\big\|\Phi^{(g)}(z_{s}^{x})-\Phi^{(g)}(z_{s}^{y})\big\|_{l^{\infty}}ds\leq\big\|x-y\big\|_{l^{\infty}}+L^{\Phi^{(g)}}\int_{0}^{t}\big\|z_{s}^{x}-z_{s}^{y}\big\|_{l^{\infty}}ds$
@@
$\|z_{t}^{x}-z_{t}^{y}\|_{l^{\infty}}\leq\left(\|x-y\|_{l^{\infty}}\right)e^{L^{\Phi^{(g)}}t}$
@@
$\|\Psi^{(g)}(x)-\Psi^{(g)}(y)\|_{l^{\infty}}=\|z_{1}^{x}-z_{1}^{y}\|_{l^{\infty}}\leq\|x-y\|_{l^{\infty}}e^{L^{\Phi^{(g)}}}$ (21)
Define $\Psi\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\bigcirc_{g=1}^{G}\Psi^{(g)}$. We try to show that the Lipschitz constant $L^{\Psi}$ is at-most $\prod_{g=1}^{G}e^{L^{\Phi^{(g)}}}$ using induction. The base case is already proved in inequality (21); for the step of the induction, consider the induction assumption bellow :
$\big\|\bigcirc_{g=1}^{G-1}\,\Psi^{(g)}(x)-\bigcirc_{g=1}^{G-1}\,\Psi^{(g)}(y)\big\|\leq\small\prod_{g=1}^{G-1}e^{L^{{\Phi^{{(g)}}}}}\big\|x-y\big\|$
@@
$\big\|\bigcirc_{g=1}^{G}\,\Psi^{(g)}(x)-\bigcirc_{g=1}^{G}\,\Psi^{(g)}(y)\big\|\leq e^{L^{\Phi^{(G)}}}\big\|\bigcirc_{g=1}^{G-1}\,\Psi^{(g)}(x)-\bigcirc_{g=1}^{G-1}\,\Psi^{(g)}(y)\big\|\leq\small\prod_{g=1}^{G}e^{L^{\Phi^{(g)}}}\big\|x-y\big\|$ (22)
So one can deduce the Lipschitz constant is at-most $\small\prod_{g=1}^{G}e^{L^{\Phi^{(g)}}}$.
@@
$\big\|\bigcirc_{g=1}^{G-1}\varphi^{(g)}(x)-\bigcirc_{g=1}^{G-1}\,\Psi^{(g)}(x)\big\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}\leq\sum_{g=1}^{G-1}\left(2\|\omega^{(g)}(\frac{d}{2n})\|_{l^{\infty}}\prod_{j=g}^{G-1}e^{L^{V^{(j)}}}\right)$
@@
$\begin{split}\big\|\varphi(x)-\bigcirc_{g=1}^{G}\,\Psi^{(g)}(x)\big\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}&=\big\|\bigcirc_{g=1}^{G}\varphi^{(g)}(x)-\bigcirc_{g=1}^{G}\,\Psi^{(g)}(x)\big\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}\\ &\leq\big\|\bigcirc_{g=1}^{G}\,\varphi^{(g)}(x)-\varphi^{(G)}\circ\bigcirc_{g=1}^{G-1}\Psi^{(g)}(x)\big\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}\\ &\quad\quad+\big\|\varphi^{(G)}\circ\bigcirc_{g=1}^{G-1}\Psi^{(g)}(x)-\bigcirc_{g=1}^{G}\,\Psi^{(g)}(x)\big\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}\\ &\leq\big\|\bigcirc_{g=1}^{G-1}\varphi^{(g)}(x)-\bigcirc_{g=1}^{G-1}\,\Psi^{(g)}(x)\big\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}e^{L^{V^{(G)}}}+2\|\omega^{(G)}(\frac{d}{2n})\|_{l^{\infty}}e^{L^{V^{(G)}}}\end{split}$
@@
$\big\|\varphi(x)-\bigcirc_{g=1}^{G}\,\Psi^{(g)}(x)\big\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}\leq\sum_{g=1}^{G}\left(2\|\omega^{(g)}(\frac{d}{2n})\|_{l^{\infty}}\prod_{j=g}^{G}e^{L^{V^{(j)}}}\right)$
@@ B.1.2 Differentiable Case
Let $\varphi\in\operatorname{\mathcal{H}}_{d}^{G,s}([0,1]^{d})$ and $1\leq s$; then, there exists ReLU neural ODEs $\{\Psi^{(g)}\}_{g=1}^{G}\subset\mathcal{X}^{0}$ such that $\Psi\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\bigcirc_{g=1}^{G}\,\Psi^{(g)}$ satisfies the approximation guarantee:
$\begin{split}\big\|\varphi-\Psi\big\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}\leq\sum_{g=1}^{G}\left(2\|\omega^{(g)}(N_{1},\dots,N_{d},L)\|_{l^{\infty}}\prod_{j=g}^{G}e^{L^{V^{(j)}}}\right)\end{split}$ (28)
Where $(\omega^{(g)})_{j}=85(s+1)^{d}8^{s}\|V^{(g)}_{j}\|_{C^{s}([0,1]^{d})}({N_{j}}{L})^{-2s/d}$ each $\omega^{(g)}$. The right hand side goes to zero as $L\rightarrow\infty$ and $N_{j}\rightarrow\infty$ for each $j\in[G]$.
Moreover, $\Psi$ is a compactly-supported $C^{s}$-diffeomorphism t on $\mathbb{R}^{d}$ supported on $(0,1)^{d}$, with Lipschitz constant at-most $\prod_{g=1}^{G}e^{L^{\Psi^{(g)}}}$ and $\Psi^{(g)}$ is a ReLU MLP of width less than or equal to $\sum_{j=1}^{d}17s^{d+1}3^{d}d(N^{g}_{j}+2)\log_{2}(8N^{g}_{j})$ and depth $18s^{2}(L^{g}+2)\log_{2}(4L^{g})+2(d+1)$ where $N^{g}_{j},L^{g}\in\mathbb{N}^{+}$ for each $1\leq j\leq d,$ and $1\leq g\leq G$.
@@
Proof : As $\varphi\in\operatorname{\mathcal{H}}_{d}^{G,s}([0,1]^{d})$ in (7) define $\varphi_{g}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\operatorname{Flow}{}(V_{g})$ for every $g\in[G]$. So one can write:
$\varphi=\bigcirc_{g=1}^{G}\,\varphi^{(g)}(x)$ (29)
We use Lemma 20 for each $\varphi^{(g)}$ to find Neural ODEs $\{\Psi^{(g)}\}_{g=1}^{G}$. One can write $\Psi^{(g)}(x)=z_{1}^{x}$ as in (23), then as $\Phi^{(g)}$ is Lipschitz we can write:
$\big\|z_{t}^{x}-z_{t}^{y}\big\|_{l^{\infty}}\leq\big\|x-y\big\|_{l^{\infty}}+\int_{0}^{t}\big\|\Phi^{(g)}(z_{s}^{x})-\Phi^{(g)}(z_{s}^{y})\big\|_{l^{\infty}}ds\leq\big\|x-y\big\|_{l^{\infty}}+L^{\Phi}\int_{0}^{t}\big\|z_{s}^{x}-z_{s}^{y}\big\|_{l^{\infty}}ds$
@@
$\|z_{t}^{x}-z_{t}^{y}\|_{l^{\infty}}\leq\left(\|x-y\|_{l^{\infty}}\right)e^{L^{\Phi^{{(g)}}}t}$
@@
$\big\|\Psi^{(g)}(x)-\Psi^{(g)}(y)\big\|_{l^{\infty}}=\big\|z_{1}^{x}-z_{1}^{y}\big\|_{l^{\infty}}\leq\left(\|x-y\|_{l^{\infty}}\right)e^{L^{\Phi^{{(g)}}}}$ (30)
Define $\Psi\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\bigcirc_{g=1}^{G}\Psi^{(g)}$. We try to show that the Lipschitz constant $L^{\Psi}$ is at-most $\prod_{g=1}^{G}e^{L^{\Phi^{(g)}}}$ using induction. The base case is already proved in inequality (30); for the step of the induction, consider the induction assumption bellow :
$\big\|\bigcirc_{g=1}^{G-1}\,\Psi^{(g)}(x)-\bigcirc_{g=1}^{G-1}\,\Psi^{(g)}(y)\big\|_{l^{\infty}}\leq\small\prod_{g=1}^{G-1}e^{L^{\Phi^{(g)}}}\big\|x-y\big\|_{l^{\infty}}$
@@
$\big\|\bigcirc_{g=1}^{G}\,\Psi^{(g)}(x)-\bigcirc_{g=1}^{G}\,\Psi^{(g)}(y)\big\|_{l^{\infty}}\leq e^{L^{\Phi^{(G)}}}\big\|\bigcirc_{g=1}^{G-1}\,\Psi^{(g)}(x)-\bigcirc_{g=1}^{G-1}\,\Psi^{(g)}(y)\big\|_{l^{\infty}}\leq\small\prod_{g=1}^{G}e^{L^{\Phi^{(g)}}}\big\|x-y\big\|_{l^{\infty}}$ (31)
So one can deduce the Lipschitz constant is at-most $\small\prod_{g=1}^{G}e^{L^{\Phi^{(g)}}}$.
Note that one can apply the same procedure of (30) for $\varphi^{(g)}$ and conclude
$L^{\varphi^{(g)}}=e^{L^{V^{(g)}}}$ (32)
@@
$\big\|\bigcirc_{g=1}^{G-1}\varphi^{(g)}(x)-\bigcirc_{g=1}^{G-1}\,\Psi^{(g)}(x)\big\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}\leq\sum_{g=1}^{G-1}\left(2\|\omega^{(g)}(N_{1},\dots,N_{d},L)\|_{l^{\infty}}\prod_{j=g}^{G-1}e^{L^{V^{(j)}}}\right)$

### v2

@@ 5 Conclusion
In this article, we have shown that incremental generation is both necessary and sufficient for universal flow-based approximation, while non-incremental flows form a non-universal subset. By coupling dynamical, algebraic, and approximation-theoretic tools, we established quantitative rates and demonstrated that lifted incremental flows yield structured universality for both functions and probability measures.
@@
The authors would like to thank Paulo Varandas for his encouraging suggestions and for verifying our arguments in proving the meagreness of the set of flows and to Giulia Livieris for her incredibly helpful feedback during the final stages of the manuscript. The authors would also like to thank Andrew Warren for the helpful references.
A. Kratsios and H. Rouhvarzi would like to acknowledge financial support from an NSERC Discovery Grant No. RGPIN-2023-04482 and No. DGECR-2023-00230. They also acknowledge that resources used in preparing this research were provided, in part, by the Province of Ontario, the Government of Canada through CIFAR, and companies sponsoring the Vector Institute [^5] https://vectorinstitute.ai/partnerships/current-partners/.
@@ Appendix A Proof of the Negative Result
Then, given any neighbourhood $U$ of the image of $\gamma$, there exists a flow $\psi$ on $M$ satisfying:
@@
Also, Lemma 17 property (a) makes the perturbed function in Lemma 18 to have support on a finite number of neighborhoods of the curves joining $p_{i}$ to $q_{i}$. As $h$ is the result of Lemma 18 take $\eta$ and those neighbourhoods small enough such that these neighbourhoods stay inside $P$. This makes the function $h$ to be compactly supported on $P$. As a result, $g$ must be compactly supported on $P$.
@@
As $D\text{exp}_{x}(0)=Id$ there exists $L>1$ tending to 1 as $r\rightarrow 0$ s.t. :
$d(f_{1}(z),f(z))\leq L\|(F-R)\circ\mathrm{exp}_{x}^{-1}(z)\|\leq L\|F-R\|r\leq L\|F-R\|\varepsilon$ (16)
@@
For every $C^{0}$-Perturbation $g$ of $f_{2}$ in $\operatorname{\mathcal{H}}_{d}(P)$ one has: $0\neq\text{deg}(f_{2}^{k}(x)-x,0)=\text{deg}(g^{k}(x)-x,0)$. There exists a point $z$ s.t. $g^{k}(z)-z=0$ so $g^{k}$ has a fixed point. Thus, $g$ has a periodic point of period $k>1$ that is not contained in a continuum subset of the set of $k$-periodic points. It means $f_{2}$ has an open neighbourhood. This proves the openness.
Proof of Theorem 2: Since every neural ODE is, by definition, a flowable homeomorphisms supported on any such $\delta,1-\delta]^{d}$ then the set of neural ODEs we mentioned is itself meagre in $\operatorname{\mathcal{H}}_{d}([0,1]^{d})$.
Proof of Theorem 1: Suppose by contradiction that for every $\varphi\in\operatorname{\mathcal{H}}_{d}([0,1]^{d})$ there exists $\varphi_{\sigma}\in\mathrm{NODE}_{\sigma}([0,1]^{d})$ such that $\|\varphi_{\sigma}-\varphi\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}<\frac{\epsilon}{2}$. Then there exists $\delta>0$ and a bump function $\psi$ which is one on $\delta,1-\delta]^{d}$ and zero out of $\frac{\delta}{2},1-\frac{\delta}{2}]^{d}$ and $\|\varphi_{\sigma}\circ\psi-\varphi_{\sigma}\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}<\frac{\epsilon}{2}$; then, $\|\varphi_{\sigma}\circ\psi-\varphi\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}<\epsilon$. As $\epsilon$ is arbitrary and $\varphi_{\sigma}\circ\psi\in\mathrm{NODE}_{\sigma}((0,1)^{d})$ this contradicts with Theorem 2.
@@ B.1.1 General Case
Proof of Proposition 7: Consider vector fields $\tilde{V}(x,y)=(-\pi(y-\frac{1}{2}),\pi(x-\frac{1}{2}))$ and $\tilde{W}(x,y)=(-x+\frac{1}{2},0)$ which are $\pi$ rotation around the point $p=(\frac{1}{2},\frac{1}{2})$ and squeeze to the line $x=\frac{1}{2}$ (call this line $l$). Consider the bump function $\rho$ which is $1$ on the circle with radius $r=\frac{1}{8}$ around $p$ and zero out of the circle with radius $R=\frac{1}{4}$. $V=\tilde{V}\circ\rho,W=\tilde{W}\circ\rho$ are compactly supported on $0,1]^{d}$. Denote time one map of $V,W$ as $\varphi_{V},\varphi_{W}$. Then $\varphi_{V}\circ\varphi_{W}\notin\mathrm{Flow}$ . suppose not, i.e. there exists a vector field $F$ which its time one flow $\varphi_{F}$ is equal to $\varphi_{V}\circ\varphi_{W}$.
Note that for the composition $\varphi_{V}\circ\varphi_{W}$ any point $a\in B_{r}(p)\setminus l$ converges to zero on iteration but any $b\in l$ has periodic point two. Consider the trajectory $\gamma$ of the flow of $\varphi_{V}\circ\varphi_{W}$. For any point $c=r_{c}e^{i\theta}$ (considering polar coordinate) on a neighbourhood $S$ of the point $\varphi_{V}\circ\varphi_{W}(b)$ on $\gamma$, radius of $\varphi_{V}\circ\varphi_{W}(c)$ will be smaller than $r_{c}$ as a result radius of any point other than $b$ on $(\varphi_{V}\circ\varphi_{W})^{2}(S)$ will be smaller but $(\varphi_{V}\circ\varphi_{W})^{2}(b)=b$ so there will be a discontinuity at the point $b$ but non-fixed periodic point of time one flow has to lie on an invariant embedded circle but the only invariant set is the line $l$, which is a contradiction.
@@
Proof of  Theorem 8: As $\varphi\in\operatorname{\mathcal{H}}_{d}^{T,s}([0,1]^{d})$ in (7) define $\varphi_{t}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\operatorname{Flow}{}(V_{t})$ for every $t\in[T]$. So one can write:
$\varphi=\bigcirc_{t=1}^{T}\,\varphi^{(t)}(x)$ (20)
We use Lemma 20 for each $\varphi^{(t)}$ to find Neural ODEs $\{\Psi^{(t)}\}_{t=1}^{T}$. One can write $\Psi^{(t)}(x)=z_{1}^{x}$ as in (17), then as $\Phi^{(t)}$ is Lipschitz we can write:
$\big\|z_{t}^{x}-z_{t}^{y}\big\|_{l^{\infty}}\leq\big\|x-y\big\|_{l^{\infty}}+\int_{0}^{t}\big\|\Phi^{(t)}(z_{s}^{x})-\Phi^{(t)}(z_{s}^{y})\big\|_{l^{\infty}}ds\leq\big\|x-y\big\|_{l^{\infty}}+L^{\Phi^{(t)}}\int_{0}^{t}\big\|z_{s}^{x}-z_{s}^{y}\big\|_{l^{\infty}}ds$
@@
$\|z_{t}^{x}-z_{t}^{y}\|_{l^{\infty}}\leq\left(\|x-y\|_{l^{\infty}}\right)e^{L^{\Phi^{(t)}}t}$
@@
$\|\Psi^{(t)}(x)-\Psi^{(t)}(y)\|_{l^{\infty}}=\|z_{1}^{x}-z_{1}^{y}\|_{l^{\infty}}\leq\|x-y\|_{l^{\infty}}e^{L^{\Phi^{(t)}}}$ (21)
Define $\Psi\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\bigcirc_{t=1}^{T}\Psi^{(t)}$. We try to show that the Lipschitz constant $L^{\Psi}$ is at-most $\prod_{t=1}^{T}e^{L^{\Phi^{(t)}}}$ using induction. The base case is already proved in inequality (21); for the step of the induction, consider the induction assumption bellow :
$\big\|\bigcirc_{t=1}^{T-1}\,\Psi^{(t)}(x)-\bigcirc_{t=1}^{T-1}\,\Psi^{(t)}(y)\big\|\leq\small\prod_{t=1}^{T-1}e^{L^{{\Phi^{{(t)}}}}}\big\|x-y\big\|$
@@
$\big\|\bigcirc_{t=1}^{T}\,\Psi^{(t)}(x)-\bigcirc_{t=1}^{T}\,\Psi^{(t)}(y)\big\|\leq e^{L^{\Phi^{(T)}}}\big\|\bigcirc_{t=1}^{T-1}\,\Psi^{(t)}(x)-\bigcirc_{t=1}^{T-1}\,\Psi^{(t)}(y)\big\|\leq\small\prod_{t=1}^{T}e^{L^{\Phi^{(t)}}}\big\|x-y\big\|$ (22)
So one can deduce the Lipschitz constant is at-most $\small\prod_{t=1}^{T}e^{L^{\Phi^{(t)}}}$.
@@
$\big\|\bigcirc_{t=1}^{T-1}\varphi^{(t)}(x)-\bigcirc_{t=1}^{T-1}\,\Psi^{(t)}(x)\big\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}\leq\sum_{t=1}^{T-1}\left(2\|\omega^{(t)}(\frac{d}{2n})\|_{l^{\infty}}\prod_{j=t}^{T-1}e^{L^{V^{(j)}}}\right)$
@@
$\begin{split}\big\|\varphi(x)-\bigcirc_{t=1}^{T}\,\Psi^{(t)}(x)\big\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}&=\big\|\bigcirc_{t=1}^{T}\varphi^{(t)}(x)-\bigcirc_{t=1}^{T}\,\Psi^{(t)}(x)\big\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}\\ &\leq\big\|\bigcirc_{t=1}^{T}\,\varphi^{(t)}(x)-\varphi^{(T)}\circ\bigcirc_{t=1}^{T-1}\Psi^{(t)}(x)\big\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}\\ &\quad\quad+\big\|\varphi^{(T)}\circ\bigcirc_{t=1}^{T-1}\Psi^{(t)}(x)-\bigcirc_{t=1}^{T}\,\Psi^{(t)}(x)\big\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}\\ &\leq\big\|\bigcirc_{t=1}^{T-1}\varphi^{(t)}(x)-\bigcirc_{t=1}^{T-1}\,\Psi^{(t)}(x)\big\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}e^{L^{V^{(T)}}}+2\|\omega^{(T)}(\frac{d}{2n})\|_{l^{\infty}}e^{L^{V^{(T)}}}\end{split}$
@@
$\big\|\varphi(x)-\bigcirc_{t=1}^{T}\,\Psi^{(t)}(x)\big\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}\leq\sum_{t=1}^{T}\left(2\|\omega^{(t)}(\frac{d}{2n})\|_{l^{\infty}}\prod_{j=t}^{T}e^{L^{V^{(j)}}}\right)$
@@ B.1.2 Differentiable Case
Let $\varphi\in\operatorname{\mathcal{H}}_{d}^{T,s}([0,1]^{d})$ and $1\leq s$; then, there exists ReLU neural ODEs $\{\Psi^{(t)}\}_{t=1}^{T}\subset\mathcal{X}^{0}$ such that $\Psi\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\bigcirc_{t=1}^{T}\,\Psi^{(t)}$ satisfies the approximation guarantee:
$\begin{split}\big\|\varphi-\Psi\big\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}\leq\sum_{t=1}^{T}\left(2\|\omega^{(t)}(N_{1},\dots,N_{d},L)\|_{l^{\infty}}\prod_{j=t}^{T}e^{L^{V^{(j)}}}\right)\end{split}$ (28)
Where $(\omega^{(t)})_{j}=85(s+1)^{d}8^{s}\|V^{(t)}_{j}\|_{C^{s}([0,1]^{d})}({N_{j}}{L})^{-2s/d}$ each $\omega^{(t)}$. The right hand side goes to zero as $L\rightarrow\infty$ and $N_{j}\rightarrow\infty$ for each $j\in[T]$.
Moreover, $\Psi$ is a compactly-supported homeomorphism t on $\mathbb{R}^{d}$ supported on $(0,1)^{d}$, with Lipschitz constant at-most $\prod_{t=1}^{T}e^{L^{\Psi^{(t)}}}$ and $\Psi^{(t)}$ is a ReLU MLP of width less than or equal to $\sum_{j=1}^{d}17s^{d+1}3^{d}d(N^{t}_{j}+2)\log_{2}(8N^{t}_{j})$ and depth $18s^{2}(L^{t}+2)\log_{2}(4L^{t})+2(d+1)$ where $N^{t}_{j},L^{t}\in\mathbb{N}^{+}$ for each $1\leq j\leq d,$ and $1\leq t\leq T$.
@@
Proof : As $\varphi\in\operatorname{\mathcal{H}}_{d}^{T,s}([0,1]^{d})$ in (7) define $\varphi_{t}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\operatorname{Flow}{}(V_{t})$ for every $t\in[T]$. So one can write:
$\varphi=\bigcirc_{t=1}^{T}\,\varphi^{(t)}(x)$ (29)
We use Lemma 21 for each $\varphi^{(t)}$ to find Neural ODEs $\{\Psi^{(t)}\}_{t=1}^{T}$. One can write $\Psi^{(t)}(x)=z_{1}^{x}$ as in (23), then as $\Phi^{(t)}$ is Lipschitz we can write:
$\big\|z_{t}^{x}-z_{t}^{y}\big\|_{l^{\infty}}\leq\big\|x-y\big\|_{l^{\infty}}+\int_{0}^{t}\big\|\Phi^{(t)}(z_{s}^{x})-\Phi^{(t)}(z_{s}^{y})\big\|_{l^{\infty}}ds\leq\big\|x-y\big\|_{l^{\infty}}+L^{\Phi^{(t)}}\int_{0}^{t}\big\|z_{s}^{x}-z_{s}^{y}\big\|_{l^{\infty}}ds$
@@
$\|z_{t}^{x}-z_{t}^{y}\|_{l^{\infty}}\leq\left(\|x-y\|_{l^{\infty}}\right)e^{L^{\Phi^{{(t)}}}t}$
@@
$\big\|\Psi^{(t)}(x)-\Psi^{(t)}(y)\big\|_{l^{\infty}}=\big\|z_{1}^{x}-z_{1}^{y}\big\|_{l^{\infty}}\leq\left(\|x-y\|_{l^{\infty}}\right)e^{L^{\Phi^{{(t)}}}}$ (30)
Define $\Psi\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\bigcirc_{t=1}^{T}\Psi^{(t)}$. We try to show that the Lipschitz constant $L^{\Psi}$ is at-most $\prod_{t=1}^{T}e^{L^{\Phi^{(t)}}}$ using induction. The base case is already proved in inequality (30); for the step of the induction, consider the induction assumption bellow :
$\big\|\bigcirc_{t=1}^{T-1}\,\Psi^{(t)}(x)-\bigcirc_{t=1}^{T-1}\,\Psi^{(t)}(y)\big\|_{l^{\infty}}\leq\small\prod_{t=1}^{T-1}e^{L^{\Phi^{(t)}}}\big\|x-y\big\|_{l^{\infty}}$
@@
$\big\|\bigcirc_{t=1}^{T}\,\Psi^{(t)}(x)-\bigcirc_{t=1}^{T}\,\Psi^{(t)}(y)\big\|_{l^{\infty}}\leq e^{L^{\Phi^{(T)}}}\big\|\bigcirc_{t=1}^{T-1}\,\Psi^{(t)}(x)-\bigcirc_{t=1}^{T-1}\,\Psi^{(t)}(y)\big\|_{l^{\infty}}\leq\small\prod_{t=1}^{T}e^{L^{\Phi^{(t)}}}\big\|x-y\big\|_{l^{\infty}}$ (31)
So one can deduce the Lipschitz constant is at-most $\small\prod_{t=1}^{T}e^{L^{\Phi^{(t)}}}$.
Note that one can apply the same procedure of (30) for $\varphi^{(t)}$ and conclude
$L^{\varphi^{(t)}}=e^{L^{V^{(t)}}}$ (32)
@@
$\big\|\bigcirc_{t=1}^{T-1}\varphi^{(t)}(x)-\bigcirc_{t=1}^{T-1}\,\Psi^{(t)}(x)\big\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}\leq\sum_{t=1}^{T-1}\left(2\|\omega^{(t)}(N_{1},\dots,N_{d},L)\|_{l^{\infty}}\prod_{j=t}^{T-1}e^{L^{V^{(j)}}}\right)$